OpenStudy (anonymous):

Find an equation of the line that has slope m and y-intercept b. (Let x be the independent variable and y be the dependent variable.) m = −9/-2; b = 3/4

OpenStudy (anonymous):

Equation of a line in slope-intercept form is y=mx+b.

OpenStudy (anonymous):

clueless here

OpenStudy (anonymous):

Just sub in the values they gave you.

Directrix (directrix):

The slope-intercept form for a line is given above by pjs. The intercept refers to the y-intercept. y = (-9/-2) x + 3/4

Directrix (directrix):

Question: Is the slope (-9/-2) or is that a typo?

OpenStudy (anonymous):

the whole thing is negative but i didn't wan to put the negative on only one number and then get a different answer.

Directrix (directrix):

Hey,( -9/-2) is not the same as - 9/2.

OpenStudy (anonymous):

yea but the whole fraction is not negative.. not just the 9

Directrix (directrix):

Example: - (1/2) = (-1)/2 = (1/(-2) but NOT (-1)/(-2) because (-1)/(-2) = + 1/2

OpenStudy (anonymous):

your right. my bad on that one lol.. didn't quite know how to place it without paper and pen. but your answer was correct

Directrix (directrix):

I have to redo the answer.

OpenStudy (anonymous):

no it was right!

Directrix (directrix):

y = (-9/2) x + 3/4

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

do you get this one : Question Part Points Submissions Used Find an equation of the line that passes through the point and has the indicated slope m. (Let x be the independent variable and y be the dependent variable.) (5, −3); m = −1